# Critical point representation of the mutual information in the sparse stochastic block model

### v1

@@ 2.1 Uniform Lipschitz continuity of the free energy
For any $\mu\in\mathcal{M}_{+}$, we have $\overline{F}_{N}(t,\mu)=\widetilde{F}_{N}\big{(}t,\mu[-1,1],\overline{\mu}\big{)}$. Indeed, this is how the enriched free energy (1.22) was defined in [18]. The Lipschitz continuity of the extended free energy (2.17) will be obtained by combining the mean value theorem with the derivative expressions (2.11)-(2.12). It will also be convenient to remember from [18, Lemma 2.3] that
@@
$\big{\lvert}\partial_{x}D_{\mu}\overline{F}_{N}(t,\mu,x)\big{\rvert}\leq c\big{(}1+\lvert\log(2c)\rvert+\lvert\log(c-\lvert\Delta\rvert)\rvert\big{)}$ (2.22)
established for any $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ in [18, Lemma 3.3] imply that
$\lVert f_{t,s,u}\rVert_{\text{Lip}}\leq c\big{(}1+\lvert\log(2c)\rvert+\lvert\log(c-\lvert\Delta\rvert)\rvert\big{)}.$
@@ 2.2 Almost everywhere differentiability of the limit free energy
$\widetilde{h}^{\prime}(t,s,q+\epsilon q^{\prime})=\widetilde{h}^{\prime}(t,s,q)+\epsilon\int_{-1}^{1}\partial_{q}\widetilde{h}^{\prime}(t,s,q,u)q^{\prime}(u)\,\mathrm{d}u+o(\epsilon).$ (2.26)
@@
For simplicity of notation, we will drop the time variable and prove that whenever the function $\widetilde{h}^{\prime}:\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]\to\mathbb{R}$ is Gateaux differentiable jointly in its two variables at the point $(s,q)\in\mathbb{R}_{>0}\times\mathcal{D}[0,1]$, then the function $h:\mathcal{M}_{+}\to\mathbb{R}$ used to define it according to (2.24) without the time variable is Gateaux differentiable at the point $s\mathcal{F}^{-1}(q)$. Generalizing the proof to incorporate the time variable requires only minor changes in notation. Let $\mu:=s\mathcal{F}^{-1}(q)$, and fix $\nu\in\mathrm{Adm}(\mathcal{M}_{+},\mu)$ as well as $\epsilon>0$ small enough so that $\mu+\epsilon\nu\in\mathcal{M}_{+}$. We define $s^{\prime}:=\nu[-1,1]$, and observe that by Gateaux differentiability of $\widetilde{h}^{\prime}$ at $(s,q)\in\mathbb{R}_{>0}\times\mathcal{D}[0,1]$,
$h(\mu+\epsilon\nu)-h(\mu)$ $=\widetilde{h}^{\prime}\bigg{(}s+\epsilon s^{\prime},q+\epsilon\frac{s^{\prime}F_{\overline{\nu}-\overline{\mu}}}{s+\epsilon s^{\prime}}\bigg{)}-\widetilde{h}^{\prime}(s,q)$
$=\epsilon\bigg{(}s^{\prime}\partial_{s}\widetilde{h}^{\prime}(s,q)+\frac{s^{\prime}}{s}\int_{-1}^{1}\partial_{q}\widetilde{h}^{\prime}(s,q,u)F_{\overline{\nu}-\overline{\mu}}(u)\,\mathrm{d}u\bigg{)}+o(\epsilon).$
@@
$h(\mu+\epsilon\nu)-h(\mu)=\epsilon\int_{-1}^{1}D_{\mu}(w)\,\mathrm{d}\nu(w)+o(\epsilon)$
@@
[…]
@@
If $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]\to\mathbb{R}$ is locally Lipschitz continuous in its first two variables and uniformly Lipschitz continuous with respect to the $L^{2}$-norm in its third variable, then it is Gateaux differentiable jointly in its three variables (in the sense of (2.26)) on a subset of $\mathbb{R}_{>0}\times\mathbb{R}_{>0}\times\mathcal{R}[0,1]$ that is dense in $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]$.
@@
For simplicity of notation, we will instead prove that a function $h:\mathcal{D}[0,1]\to\mathbb{R}$ that is locally Lipschitz continuous with respect to the $L^{2}$-norm is Gateaux differentiable on a subset of $\mathcal{R}[0,1]$ that is dense in $\mathcal{D}[0,1]$. Our proof generalizes with only modifications in notation to the setting of a function defined on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]$. Up to multiplying the function $h$ with a cutoff function that vanishes when the total mass of the measure becomes larger than some predetermined threshold, we will also assume without loss of generality that $h:\mathcal{D}[0,1]\to\mathbb{R}$ is in fact uniformly Lipschitz continuous with respect to the $L^{2}$-norm. The proof proceeds in three steps. First, we construct a Lipschitz extension $\overline{h}:L^{2}[-1,1]\to\mathbb{R}$ of $h$, then we show that there exists a Gaussian null set $\mathcal{N}$ such that $h$ is Gateaux differentiable on $\mathcal{D}[0,1]\setminus\mathcal{N}$, and finally we prove that $\mathcal{R}[0,1]\setminus\mathcal{N}$ is dense in $\mathcal{D}[0,1]$.
Step 1: constructing the Lipschitz extension. We denote by $P:L^{2}[-1,1]\to\mathcal{D}[0,1]$ the orthogonal projection onto the closed convex set $\mathcal{D}[0,1]$. We refer the reader to [19, Exercise 2.16] and its solution for a justification that this projection $P$ is well-defined based on the fact that $\mathcal{D}[0,1]$ is a closed convex subset of $L^{2}[-1,1]$. It is also shown there using the convexity of $\mathcal{D}[0,1]$ that, for every $q\in L^{2}[-1,1]$ and $p\in\mathcal{D}[0,1]$,
@@
Step 3: density of $\mathcal{R}[0,1]\setminus\mathcal{N}$ in $\mathcal{D}[0,1]$. The set $\mathcal{R}[0,1]$ has dense linear span in $L^{2}[-1,1]$, so the separability of $L^{2}[-1,1]$ allows us to find a sequence $(w_{n})_{n\geq 1}\subset\mathcal{R}[0,1]$ whose linear span is dense in $L^{2}[-1,1]$. Re-scaling each $w_{n}$, we can assume without loss of generality that $\lvert w_{n}\rvert_{L^{2}}\leq 1$ and that $\lim_{n\to+\infty}\lvert w_{n}\rvert_{L^{2}}=0$. Let $K$ be as in Lemma 2.6, and observe that $K\subset\mathcal{R}[0,1]$ by convexity of $\mathcal{R}[0,1]$. At this point, fix $q\in\mathcal{R}[0,1]$ as well as $\epsilon>0$. Since $q+\epsilon K$ is not a Gaussian null set, we have that $q+\epsilon K\not\subset\mathcal{N}$. In particular, there exists $y\in q+\epsilon K\setminus\mathcal{N}$. Notice that for any $\kappa\in K$, we have $(q+\epsilon\kappa)^{\prime}>0$ so in fact $y\in\mathcal{R}\setminus\mathcal{N}$ with $\lvert y-q\rvert_{L^{2}}\leq\epsilon$. This establishes the density of $\mathcal{R}[0,1]\setminus\mathcal{N}$ in $\mathcal{R}[0,1]$. Together with the density of $\mathcal{R}[0,1]$ in $\mathcal{D}[0,1]$ which can be obtained through mollifiers, this completes the proof. ∎
@@
:=$\mathcal{B}:=\big{\{}\big{(}t,s\mathcal{F}^{-1}(q)\big{)}\mid(t,s,q)\in\mathcal{A}\big{\}}\subset\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+}).$ (2.28)
@@ 2.3 Convergence of the derivative free energy
For all $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ and $\nu\in\Pr[-1,1]$, the map $r\mapsto\overline{F}_{N}(t,\mu+r\nu)$ is convex.
@@
The superposition principle for Poisson point processes implies that, jointly in $\sigma\in\Sigma_{N}$,
@@
Together with [19, Proposition 5.12] on Poisson averages, this implies that
@@
A direct computation leveraging the product rule shows that
@@
$\widetilde{\mathbf{G}}_{N}^{\prime}$ :=$:=\big{\{}\widetilde{G}_{m,k}\mid m\neq i,\,\,\,m\neq j\text{ and}k\leq\Pi_{m,r}\big{\}}\quad\text{and}\quad Z:=\big{(}(\widetilde{G}_{i,\ell})_{\ell\leq L_{1}},(\widetilde{G}_{j,\ell})_{\ell\leq L_{2}},\widetilde{\mathbf{G}}_{N}^{\prime}\big{)}.$
@@
Fix $r,\eta>0$ and a probability measure $\nu\in\Pr[-1,1]\cap\mathrm{Reg}(\mathcal{M}_{+})$. We define the measure $\nu_{\eta}\in\mathrm{Reg}(\mathcal{M}_{+})$ by $\nu_{\eta}:=\eta\nu$. The convexity property of the enriched free energy in Proposition 2.9 gives the upper bound
@@
Repeating the argument with $\eta$ replaced by $-\eta$, which is possible as $\mu,\nu\in\mathrm{Reg}(\mathcal{M}_{+})$, shows that for any $\nu\in\mathrm{Reg}(\mathcal{M}_{+})$, we have $D_{t,\mu}(\nu)=D_{\mu}f(t,\mu;\nu)$. By density of $\mathrm{Reg}(\mathcal{M}_{+})$ in $\mathcal{M}_{+}$, this completes the proof. ∎
@@ 3 Multioverlap concentration through perturbation
As will be explained below, here $(\sigma_{i}^{\ell})_{i,\ell\geq 1}$ denote i.i.d. replicas from the Gibbs measure associated with the sum of the Hamiltonian (1.21) and the two perturbation Hamiltonians that we now introduce. The first perturbation Hamiltonian will ensure the concentration of the overlap $R_{1,2}$. Given a sequence $(\epsilon_{N})_{N\geq 1}$ with $\epsilon_{N}:=N^{\gamma}$ for some $-1/8<\gamma<0$ and a standard Gaussian vector $Z_{0}:=(Z_{0,1},\ldots,Z_{0,N})$ in $\mathbb{R}^{N}$, we introduce the Gaussian perturbation Hamiltonian
$H_{N}^{\mathrm{gauss}}(\sigma,\lambda_{0}):=\mathcal{H}_{0}:=\sum_{i\leq N}\big{(}\lambda_{0}\epsilon_{N}\sigma^{*}_{i}\sigma_{i}+\sqrt{\lambda_{0}\epsilon_{N}}Z_{0,i}\sigma_{i}\big{)}$ (3.3)
@@
$Y^{\mathrm{gauss}}:=\sqrt{\lambda_{0}\epsilon_{N}}\sigma^{*}+Z_{0}.$ (3.4)
@@
$1\geq\epsilon_{N}\to 0\quad\text{and}\quad N\epsilon_{N}\to+\infty.$ (3.5)
The second perturbation Hamiltonian will ensure the validity of the Franz-de Sanctis identities [8, 18] which together with the concentration of the overlap $R_{1,2}$ give the concentration of all the multioverlaps (3.2). We consider a sequence $(s_{N})_{N\geq 1}$ with $s_{N}:=N^{\eta}$ for some $14/5<\eta<$ 1, so that in particular
@@
[…]
@@
This corresponds to [18, Lemma 4.1] but we provide a proof for completeness. A direct computation reveals that
@@
$\big{\lvert}H_{N}^{\mathrm{gauss}}(\sigma,\lambda_{0})\big{\rvert}\leq N\epsilon_{N}+\sqrt{\epsilon_{N}}\sum_{i\leq N}\lvert Z_{0,i}\rvert$
@@
$\big{\lvert}\overline{F}_{N}(t,\mu,\lambda)-\overline{F}_{N}(t,\mu)\big{\rvert}\leq\epsilon_{N}+\sqrt{\epsilon_{N}}\operatorname{\mathbb{E}}Z_{0,1}+\frac{s_{N}}{N}\sum_{k\geq 1}\Big{(}\log(1+\lambda_{k})+\frac{\lambda_{k}}{1-\lambda_{k}}\Big{)}.$
@@
where $\operatorname{\mathbb{E}}_{\lambda}$ denotes the average with respect to the uniform random variables $\lambda_{k}$ on $[2^{-k-1},2^{-k}]$.
@@
and think of the perturbation parameters $\lambda:=(\lambda_{k})_{k\geq 0}$ as being sampled independently with $\lambda_{k}$ uniform on $[2^{-k-1},2^{-k}]$ for $k\geq 0$. The concentration function $\nu_{N}$ is bounded from above by a constant independent of $N$ by [18, Proposition B.4], so (3.13) follows from [8, Theorem 2.2]. This completes the proof. ∎
@@
We discuss briefly how to extend the results of the present paper and cover the more general case when the measure $P^{*}$ is arbitrary with support in $[-1,1]$ in place of $\{-1,1\}$, with the understanding that we stick with the formula (1.4) to compute the link probabilities. In this case, the description of the limit law of $(\sigma^{\ell}_{i})_{i,\ell\geq 1}$ must naturally be modified from (S1)-(S2), since these explicitly refer to ±$\pm 1$ random variables. Instead, we would first need to sample i.i.d. random probability measures $(\mu_{i})_{i\geq 1}$, each with law $\Xi\in\Pr(\Pr[-1,1])$, and then conditionally on $(\mu_{i})_{i\geq 1}$, to sample $\sigma_{i}^{\ell}$ according to the law $\mu_{i}$ independently over $i,\ell\geq 1$. One key observation though is that Lemma 3.4 remains valid as stated, provided that we understand that $\nu$ is the law of the mean value of $\mu_{1}$, that is,
@@
Moreover, the cavity computations ultimately only require us to be able to evaluate moments of the form described by Lemma 3.4. So even in this more general setting, we do not really need to keep track of the more complex object $\Xi\in\Pr(\Pr[-1,1])$, and can proceed as in the ±$\pm 1$ case by focusing on this simpler measure $\nu\in\mathcal{M}_{p}$.

### v2

@@ 2.1 Uniform Lipschitz continuity of the free energy
For any $\mu\in\mathcal{M}_{+}$, we have $\overline{F}_{N}(t,\mu)=\widetilde{F}_{N}\big(t,\mu[-1,1],\overline{\mu}\big)$, where $\overline{\mu}$ is the probability measure obtained by normalizing $\mu$ as in (2.7). Indeed, this is how the enriched free energy (1.22) was defined in [19]. The Lipschitz continuity of the extended free energy (2.18) will be obtained by combining the mean value theorem with the derivative expressions (2.12)-(2.13). It will also be convenient to remember from [19, Lemma 2.3] that
@@
[…]
established for any $(t,\mu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$ in [19, Lemma 3.3] imply that
[…]
@@ 2.2 Almost everywhere differentiability of the limit free energy
$\widetilde{h}^{\prime}(t,s,q+\varepsilon q^{\prime})=\widetilde{h}^{\prime}(t,s,q)+\varepsilon\int_{-1}^{1}\partial_{q}\widetilde{h}^{\prime}(t,s,q,u)q^{\prime}(u)\,\mathrm{d}u+o(\varepsilon).$ (2.27)
@@
For simplicity of notation, we will drop the time variable and prove that whenever the function $\widetilde{h}^{\prime}:\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]\to\mathbb{R}$ is Gateaux differentiable jointly in its two variables at the point $(s,q)\in\mathbb{R}_{>0}\times\mathcal{D}[0,1]$, then the function $h:\mathcal{M}_{+}\to\mathbb{R}$ used to define it according to (2.25) without the time variable is Gateaux differentiable at the point $s\mathcal{F}^{-1}(q)$. Generalizing the proof to incorporate the time variable requires only minor changes in notation. Let $\mu:=s\mathcal{F}^{-1}(q)$, and fix $\nu\in\mathrm{Adm}(\mathcal{M}_{+},\mu)$ as well as $\varepsilon>0$ small enough so that $\mu+\varepsilon\nu\in\mathcal{M}_{+}$. We define $s^{\prime}:=\nu[-1,1]$, and observe that by Gateaux differentiability of $\widetilde{h}^{\prime}$ at $(s,q)\in\mathbb{R}_{>0}\times\mathcal{D}[0,1]$,
$h(\mu+\varepsilon\nu)-h(\mu)$ $=\widetilde{h}^{\prime}\bigg(s+\varepsilon s^{\prime},q+\varepsilon\frac{s^{\prime}F_{\overline{\nu}-\overline{\mu}}}{s+\varepsilon s^{\prime}}\bigg)-\widetilde{h}^{\prime}(s,q)$
$=\varepsilon\bigg(s^{\prime}\partial_{s}\widetilde{h}^{\prime}(s,q)+\frac{s^{\prime}}{s}\int_{-1}^{1}\partial_{q}\widetilde{h}^{\prime}(s,q,u)F_{\overline{\nu}-\overline{\mu}}(u)\,\mathrm{d}u\bigg)+o(\varepsilon).$
@@
$h(\mu+\varepsilon\nu)-h(\mu)=\varepsilon\int_{-1}^{1}D_{\mu}(w)\,\mathrm{d}\nu(w)+o(\varepsilon)$
@@
$\mathcal{R}[0,1]:=\big\{q\in\mathcal{D}[0,1]\mid q$ $\text{ is smooth with strictly positive derivative}\big\}.$ (2.28)
@@
If $h:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]\to\mathbb{R}$ is locally Lipschitz continuous in its first two variables and locally Lipschitz continuous with respect to the $L^{2}$-norm in its third variable, then it is Gateaux differentiable jointly in its three variables (in the sense of (2.27)) on a subset of $\mathbb{R}_{>0}\times\mathbb{R}_{>0}\times\mathcal{R}[0,1]$ that is dense in $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]$.
@@
For simplicity of notation, we will instead prove that a function $h:\mathcal{D}[0,1]\to\mathbb{R}$ that is locally Lipschitz continuous with respect to the $L^{2}$-norm is Gateaux differentiable on a subset of $\mathcal{R}[0,1]$ that is dense in $\mathcal{D}[0,1]$. Our proof generalizes with only modifications in notation to the setting of a function defined on $\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\times\mathcal{D}[0,1]$. Up to multiplying the function $h$ with a cutoff function that vanishes outside of a sufficiently large ball, we will also assume without loss of generality that $h:\mathcal{D}[0,1]\to\mathbb{R}$ is in fact uniformly Lipschitz continuous with respect to the $L^{2}$-norm. The proof proceeds in three steps. First, we construct a Lipschitz extension $\overline{h}:L^{2}[-1,1]\to\mathbb{R}$ of $h$, then we show that there exists a Gaussian null set $\mathcal{N}$ such that $h$ is Gateaux differentiable on $\mathcal{D}[0,1]\setminus\mathcal{N}$, and finally we prove that $\mathcal{R}[0,1]\setminus\mathcal{N}$ is dense in $\mathcal{D}[0,1]$.
Step 1: constructing the Lipschitz extension. We denote by $P:L^{2}[-1,1]\to\mathcal{D}[0,1]$ the orthogonal projection onto the closed convex set $\mathcal{D}[0,1]$. We refer the reader to [20, Exercise 2.16] and its solution for a justification that this projection $P$ is well-defined based on the fact that $\mathcal{D}[0,1]$ is a closed convex subset of $L^{2}[-1,1]$. It is also shown there using the convexity of $\mathcal{D}[0,1]$ that, for every $q\in L^{2}[-1,1]$ and $p\in\mathcal{D}[0,1]$,
@@
Step 3: density of $\mathcal{R}[0,1]\setminus\mathcal{N}$ in $\mathcal{D}[0,1]$. The set $\mathcal{R}[0,1]$ has dense linear span in $L^{2}[-1,1]$, so the separability of $L^{2}[-1,1]$ allows us to find a sequence $(w_{n})_{n\geq 1}\subset\mathcal{R}[0,1]$ whose linear span is dense in $L^{2}[-1,1]$. Re-scaling each $w_{n}$, we can assume without loss of generality that $\lvert w_{n}\rvert_{L^{2}}\leq 1$ and that $\lim_{n\to+\infty}\lvert w_{n}\rvert_{L^{2}}=0$. Let $K$ be as in Lemma 2.6, and observe that $K\subset\mathcal{R}[0,1]$ by convexity of $\mathcal{R}[0,1]$. At this point, fix $q\in\mathcal{R}[0,1]$ as well as $\varepsilon>0$. Since $q+\varepsilon K$ is not a Gaussian null set, we have that $q+\varepsilon K\not\subset\mathcal{N}$. In particular, there exists $y\in q+\varepsilon K\setminus\mathcal{N}$. Notice that for any $\kappa\in K$, we have $(q+\varepsilon\kappa)^{\prime}>0$ so in fact $y\in\mathcal{R}\setminus\mathcal{N}$ with $\lvert y-q\rvert_{L^{2}}\leq\varepsilon$. This establishes the density of $\mathcal{R}[0,1]\setminus\mathcal{N}$ in $\mathcal{R}[0,1]$. Together with the density of $\mathcal{R}[0,1]$ in $\mathcal{D}[0,1]$ which can be obtained through mollifiers, this completes the proof. ∎
@@
$\mathcal{B}:=\big\{\big(t,s\mathcal{F}^{-1}(q)\big)\mid(t,s,q)\in\mathcal{A}\big\}\subset\mathbb{R}_{\geq 0}\times\mathrm{Reg}(\mathcal{M}_{+}).$ (2.29)
@@ 2.3 Convergence of the derivative free energy
For all $(t,\mu),(s,\nu)\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}$, the map $r\mapsto\overline{F}_{N}(t+rs,\mu+r\nu)$ is convex.
@@
The claim essentially follows from the general fact that if $S$ is a random variable and if $(X_{n})_{n\geq 1}$ are i.i.d. conditionally on $S$, then the mutual information between $S$ and $(X_{n})_{n\leq\operatorname{\operatorname{Poi}}(t)}$ is a concave function of $t$ (see e.g. the proof of [25, Proposition 4.1]). In order to lighten the notation, we only prove the result for $s=0$. (We note that for later purposes, all we really need is to apply Proposition 2.9 either with $s=0$ or with $\nu=0$; and the proof in the case $\nu=0$ really boils down to the general observation above.) Without loss of generality, we may assume that $\nu\in\Pr[-1,1]$. The superposition principle for Poisson point processes implies that, jointly in $\sigma\in\Sigma_{N}$,
@@
Together with [20, Proposition 5.12] on Poisson averages, this implies that
@@
For every $\lambda>0$ and integer $m\geq 0$, we write
$\pi(\lambda,m)=\frac{\lambda^{m}}{m!}\exp(-\lambda)$ (2.32)
to denote the probability for a Poisson random variable with parameter $\lambda$ to be equal to $m$. A direct computation leveraging the product rule shows that
@@
$\widetilde{\mathbf{G}}_{N}^{\prime}$ $:=\big\{\widetilde{G}_{m,k}\mid m\neq i,\,\,\,m\neq j\text{ and }k\leq\Pi_{m,r}\big\}\quad\text{and}\quad Z:=\big((\widetilde{G}_{i,\ell})_{\ell\leq L_{1}},(\widetilde{G}_{j,\ell})_{\ell\leq L_{2}},\widetilde{\mathbf{G}}_{N}^{\prime}\big).$
@@
The claim concerning the convergence of the derivative with respect to the $t$ variable is a direct consequence of the convexity in $t$ ensured by Proposition 2.9, see for instance [20, Proposition 2.15]. We now turn to the convergence of the derivative with respect to $\mu$. Fix $r,\eta>0$ and a probability measure $\nu\in\Pr[-1,1]\cap\mathrm{Reg}(\mathcal{M}_{+})$. We define the measure $\nu_{\eta}\in\mathrm{Reg}(\mathcal{M}_{+})$ by $\nu_{\eta}:=\eta\nu$. The convexity property of the enriched free energy in Proposition 2.9 gives the upper bound
@@
Repeating the argument with $\eta$ replaced by $\eta$, which is possible as $\mu,\nu\in\mathrm{Reg}(\mathcal{M}_{+})$, shows that for any $\nu\in\mathrm{Reg}(\mathcal{M}_{+})$, we have $D_{t,\mu}(\nu)=D_{\mu}f(t,\mu;\nu)$. By density of $\mathrm{Reg}(\mathcal{M}_{+})$ in $\mathcal{M}_{+}$, this completes the proof. ∎
@@ 3 Multioverlap concentration through perturbation
As will be explained below, here $(\sigma_{i}^{\ell})_{i,\ell\geq 1}$ denote i.i.d. replicas from the Gibbs measure associated with the sum of the Hamiltonian (1.21) and the two perturbation Hamiltonians that we now introduce. The first perturbation Hamiltonian will ensure the concentration of the overlap $R_{1,2}$. Given a sequence $(\varepsilon_{N})_{N\geq 1}$ with $\varepsilon_{N}:=N^{\gamma}$ for some $1/8<\gamma<0$ and a standard Gaussian vector $Z_{0}:=(Z_{0,1},\ldots,Z_{0,N})$ in $\mathbb{R}^{N}$, we introduce the Gaussian perturbation Hamiltonian
$H_{N}^{\mathrm{gauss}}(\sigma,\lambda_{0}):=\mathcal{H}_{0}:=\sum_{i\leq N}\big(\lambda_{0}\varepsilon_{N}\sigma^{*}_{i}\sigma_{i}+\sqrt{\lambda_{0}\varepsilon_{N}}Z_{0,i}\sigma_{i}\big)$ (3.3)
@@
$Y^{\mathrm{gauss}}:=\sqrt{\lambda_{0}\varepsilon_{N}}\sigma^{*}+Z_{0}.$ (3.4)
@@
$1\geq\varepsilon_{N}\to 0\quad\text{and}\quad N\varepsilon_{N}\to+\infty.$ (3.5)
The second perturbation Hamiltonian will ensure the validity of the Franz-de Sanctis identities [8, 19] which together with the concentration of the overlap $R_{1,2}$ give the concentration of all the multioverlaps (3.2). We consider a sequence $(s_{N})_{N\geq 1}$ with $s_{N}:=N^{\eta}$ for some $4/5<\eta<1$, so that in particular
@@
its associated free energy functionals, and by $\langle\cdot\rangle$ its corresponding Gibbs average. This means that for any bounded and measurable function $f=f(\sigma^{1},\ldots,\sigma^{n})$ of finitely many replicas,
@@
This corresponds to [19, Lemma 4.1] but we provide a proof for completeness. A direct computation reveals that
@@
$\big\lvert H_{N}^{\mathrm{gauss}}(\sigma,\lambda_{0})\big\rvert\leq N\varepsilon_{N}+\sqrt{\varepsilon_{N}}\sum_{i\leq N}\lvert Z_{0,i}\rvert$
@@
$\big\lvert\overline{F}_{N}(t,\mu,\lambda)-\overline{F}_{N}(t,\mu)\big\rvert\leq\varepsilon_{N}+\sqrt{\varepsilon_{N}}\operatorname{\mathbb{E}}Z_{0,1}+\frac{s_{N}}{N}\sum_{k\geq 1}\Big(\log(1+\lambda_{k})+\frac{\lambda_{k}}{1-\lambda_{k}}\Big).$
@@
where $\operatorname{\mathbb{E}}_{\lambda}$ denotes the average with respect to the uniform random variables $\lambda_{k}$ on $2^{-k-1},2^{-k}]$.
@@
and think of the perturbation parameters $\lambda:=(\lambda_{k})_{k\geq 0}$ as being sampled independently with $\lambda_{k}$ uniform on $2^{-k-1},2^{-k}]$ for $k\geq 0$. The concentration function $\nu_{N}$ is bounded from above by a constant independent of $N$ by [19, Proposition B.4], so (3.13) follows from [8, Theorem 2.2]. This completes the proof. ∎
@@
We discuss briefly how to extend the results of the present paper and cover the more general case when the measure $P^{*}$ is arbitrary with support in [1,1][-1,1] in place of $\{-1,1\}$, with the understanding that we stick with the formula (1.4) to compute the link probabilities. In this case, the description of the limit law of $(\sigma^{\ell}_{i})_{i,\ell\geq 1}$ must naturally be modified from (S1)-(S2), since these explicitly refer to $\pm 1$ random variables. Instead, we would first need to sample i.i.d. random probability measures $(\mu_{i})_{i\geq 1}$, each with law $\Xi\in\Pr(\Pr[-1,1])$, and then conditionally on $(\mu_{i})_{i\geq 1}$, to sample $\sigma_{i}^{\ell}$ according to the law $\mu_{i}$ independently over $i,\ell\geq 1$. One key observation though is that Lemma 3.4 remains valid as stated, provided that we understand that $\nu$ is the law of the mean value of $\mu_{1}$, that is,
@@
Moreover, the cavity computations ultimately only require us to be able to evaluate moments of the form described by Lemma 3.4. So even in this more general setting, we do not really need to keep track of the more complex object $\Xi\in\Pr(\Pr[-1,1])$, and can proceed as in the $\pm 1$ case by focusing on this simpler measure $\nu\in\mathcal{M}_{p}$.